# Diffusion Policy Policy Optimization

### v1

@@ D.1 Details of policy architectures used in all experiments
All methods from Section 5.1 use the same pre-trained policy. Fine-tuning is done using online experiences sampled from 40 parallelized MuJoCo environments [77]. Reward curves shown in Fig. 5 are evaluated by running fine-tuned policies with $\sigma^{\text{exp}}_{\text{min}}=0.001$ (i.e., without extra noise) for 40 episodes. Each episode terminates if the default conditions are met or the episode reaches 1000 timesteps. Detailed hyperparameters are listed in Table A7 and Table A8.
D.3 Descriptions of diffusion-based RL algorithm baselines in Section 5.1
@@
D.4 Additional details of DPPO implementation in all tasks
Similar to all baselines in Section D.3, we denote $N_{\theta}$ and $N_{\phi}$ the replay ratio for the actor (Diffusion Policy) and the critic (state value function) in DPPO; in practice we always set $N_{\theta}=N_{\phi}$ in DPPO, with the combined loss $\mathcal{L}=\mathcal{L}_{\theta}+\mathcal{L}_{\phi}$. Similar to usual PPO implementations [32], the batch updates in an iteration terminate when the KL divergence between $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$ reaches 1.
[…]
D.5 Additional details of Robomimic tasks and training in Section 5.2
@@ D.4 Additional details of DPPO implementation in all tasks
D.6 Descriptions of policy parameterization baselines in Section 5.2
@@
We consider unimodal Gaussian with diagonal covariance, the most commonly used policy parameterization in RL. The standard deviation for each action dimension, $\sigma_{\text{Gau}}$, is fixed during pre-training; we also tried to learn $\sigma_{\text{Gau}}$ from the dataset but we find the training very unstable. During fine-tuning $\sigma_{\text{Gau}}$ is learned starting from the same fixed value and also clipped between $0.01$ and $0.2$. Additionally we clip the sampled action to be within $3$ standard deviation from the mean. As discusses in Section D.4, we choose the PPO clipping ratio $\varepsilon$ based on the empirical clipping fraction in each task. This setup is also used in the Furniture-Bench experiments. We note that we spend significant amount of efforts tuning the Gaussian baseline, and our results with it are some of the best known ones in RL training for long-horizon manipulation tasks (exceeding our initial expectations), e.g., reaching $\sim$100success rate in Lamp with Low randomness.
@@
D.7 Additional details of Furniture-Bench tasks and training in Section 5.3
@@
To deploy state-based policies on real hardware, we utilize AprilTags [82] for part pose estimation. The Furniture-Bench [28] task suite provides AprilTags for each part and code for estimating part poses from tag detections. The process involves several steps: (1) detecting tags in the camera frame, (2) mapping tag detections to the robot frame for policy compatibility, (3) utilizing known offsets between tags and object centers in the simulator, and (4) calibrating the camera pose using an AprilTag at a known position relative to the robot base. Despite general accuracy, detections can be noisy, especially during movement or partial occlusion. To mitigate occasional large “jumps” and noise in pose estimations, we implement a simple low-pass filter (exponential moving average) and average over multiple estimates from different tags or cameras for the same object.
@@
D.8 Additional details of Avoid task from D3IL and training in Section 6
@@
D.9 Listed training hyperparameters
@@
$K^{\prime}$ 20
Actor learning rate 1e-4 1e-5 1e-5 1e-5 (decayed to 1e-6)
@@
$N_{\phi}$ 16 4 4 4
Buffer size 200000 150000 150000 150000
Batch size 256
@@
Buffer size 400000
Batch size 5000
@@
Buffer size 200000 150000 150000 150000
Batch size 256 512 512 512
DQL $\alpha_{\text{DQL}}$ 1
$N_{\theta}$ 64
$N_{\phi}$ 64
Buffer eize 400000
Batch size 5000
Table A7: Fine-tuning hyperparameters for OpenAI Gym and Robomimic tasks when comparing diffusion-based RL methods. We list hyperparameters shared by all methods first, and then method-specific ones.
Task(s) (cont’d)
Method (cont’d) Parameter (cont’d) Gym Lift, Can Square Transport
QSM $\alpha_{\text{QSM}}$ 50
$N_{\theta}$ 32
$N_{\phi}$ 32
Buffer size 200000 150000 150000 150000
Batch size 256
@@
Batch size 5000 7500 10000 10000
Table A8: Continuation of Table A7.

### v3

@@ E.1 Details of policy architectures used in all experiments
All methods from Section 5.1 use the same pre-trained policy. Fine-tuning is done using online experiences sampled from 40 parallelized MuJoCo environments [92]. Reward curves shown in Fig. 5 are evaluated by running fine-tuned policies with $\sigma^{\text{exp}}_{\text{min}}=0.001$ (i.e., without extra noise) for 40 episodes. Each episode terminates if the default conditions are met or the episode reaches 1000 timesteps. Detailed hyperparameters are listed in Table A7.
E.3 Descriptions of diffusion-based RL algorithm baselines in Section 5.1
@@
E.4 Descriptions of RL fine-tuning algorithm baselines in Section 5.2
In this subsection, we detail the baselines RLPD, Cal-QL, and IBRL. All policies $\pi_{\theta}$ are parameterized as unimodal Gaussian.
RLPD [7]:
This baseline is based on Soft Actor Critic (SAC, Haarnoja et al. [33]) — it learns an entropy-regularized state-action Q function, and then updates the actor by maximizing the Q function w.r.t. the action.
A replay buffer $\mathcal{D}$ is initialized with offline data, and online samples are added to $\mathcal{D}$. Each gradient update uses a batch of mixed 50/50 offline and online data. An ensemble of $N_{\text{critic}}$ critics is used, and at each gradient step two critics are randomly chosen. The critics are trained to minimize the Bellman residual with replay ratio $N_{\phi}$:
$\mathcal{L}_{\phi}=\mathbb{E}^{\mathcal{D}}\big{[}\|(R_{t}+\gamma_{\textsc{env}}\hat{Q}_{\phi^{\prime}}(s_{t+1},\pi_{\theta}(a_{t+1}|s_{t+1}))-\hat{Q}_{\phi}(s_{t},a_{t})\|^{2}\big{]}.$ (A14)
The target critic parameter $\phi^{\prime}$ is updated with delay. The actor minimizes the following loss with a replay ratio of 1:
$\mathcal{L}_{\theta}=\mathbb{E}^{\mathcal{D}}\big{[}-\hat{Q}_{\phi}(s_{t},a_{t})+\alpha_{\text{ent}}\log\pi_{\theta}(a_{t}|s_{t})\big{]},$ (A15)
where $\alpha_{\text{ent}}$ is the entropy coefficient (automatically tuned as in SAC starting at 1).
Cal-QL [62]:
This baseline trains the policy $\mu$ and the action-value function $Q^{\mu}$ in an offline phase and then an online phase. During the offline phase only offline data is sampled for gradient update, while during the online phase mixed 50/50 offline and online data are sampled. The critic is trained to minimize the following loss (Bellman residual and calibrated Q-learning):
ϕ=subscriptitalic-$\mathcal{L}_{\phi}=$ $\mathbb{E}^{\mathcal{D}}\big{[}\|(R_{t}+\gamma_{\textsc{env}}\hat{Q}_{\phi^{\prime}}(s_{t+1},\pi_{\theta}(a_{t+1}|s_{t+1})))-\hat{Q}_{\phi}(s_{t},a_{t})\|^{2}\big{]}$ (A16)
$+\beta_{\text{cql}}(\mathbb{E}^{\mathcal{D}}\big{[}\max(Q_{\phi}(s_{t},a_{t}),V(s_{t}))\big{]}-\mathbb{E}^{\mathcal{D}}\big{[}Q_{\phi}(s_{t},a_{t})\big{]}),$
where $\beta_{\text{cql}}$ is a weighting coefficient between Bellman residual and calibration Q-learning and $V(s_{t})$ is estimated using Monte-Carlo returns. The target critic parameter $\phi^{\prime}$ is updated with delay. The actor minimizes the following loss:
$\mathcal{L}_{\theta}=\mathbb{E}^{\mathcal{D}}\big{[}-\hat{Q}_{\phi}(s_{t},a_{t})+\alpha_{\text{ent}}\log\pi_{\theta}(a_{t}|s_{t})\big{]},$ (A17)
where $\alpha_{\text{ent}}$ is the entropy coefficient (automatically tuned as in SAC starting at 1).
IBRL [40]:
This baseline first pre-trains a policy $\mu_{\psi}$ using behavior cloning, and for fine-tuning it trains a RL policy $\pi_{\theta}$ initialized as $\mu_{\psi}$. During fine-tuning recent experiences are saved in a replay buffer $\mathcal{D}$. An ensemble of $N_{\text{critic}}$ critics is used, and at each gradient step two critics are randomly chosen. The critics are trained to minimize the Bellman residual with replay ratio $N_{\phi}$:
$\mathcal{L}_{\phi}=\mathbb{E}^{\mathcal{D}}\big{[}\|(R_{t}+\gamma_{\textsc{env}}\underset{a^{\prime}\in\{a^{IL},a^{RL}\}}{\max}\hat{Q}_{\phi^{\prime}}(s_{t+1},a^{\prime})-\hat{Q}_{\phi}(s_{t},a_{t})\|^{2}\big{]}$ (A18)
where $a^{IL}=\mu_{\psi}(s_{t+1})$ (no noise) and $a^{RL}\sim\pi_{\theta^{\prime}}(s_{t+1})$, and $\pi_{\theta^{\prime}}$ is the target actor. The target critic parameter $\phi^{\prime}$ is updated with delay. The actor minimizes the following loss with a replay ratio of 1:
$\mathcal{L}_{\theta}=-\mathbb{E}^{\mathcal{D}}\big{[}\hat{Q}_{\phi}(s_{t},a_{t})\big{]}.$ (A19)
The target actor parameter $\theta^{\prime}$ is also updated with delay.
E.5 Additional details of Franka-Kitchen tasks and training in Section 5.2
Tasks.
We consider three settings from the D4RL benchmark [29]: (1) Kitchen-Complete-v0 containing demonstrations that complete the entire task (four subtasks), (2) Kitchen-Partial-v0 containing some complete demonstrations and many ones completing only subtasks, and (3) Kitchen-Mixed-v0 containing incomplete demonstrations only.
Pre-training.
The observations and actions are normalized to $[0,1]$ using min/max statistics from the pre-training dataset. No history observation (proprioception or ground-truth object states) is used. All policies are trained with batch size 128, learning rate 1e-4 decayed to 1e-5 with a cosine schedule, and weight decay 1e-6. DPPO policies are trained with 8000 epochs. For IBRL and Cal-QL we follow the hyperparameters from the original implementations — IBRL proposes using (1) wider MLP layers and (2) dropout during pre-training, which we follow too. We use $T_{a}=4$ for DPPO; we also tried to use the same action chunk size with IBRL, RLPD, and Cal-QL, but we find for all of them $T_{a}=1$ leads to better performance.
Fine-tuning.
With DPPO, policies are fine-tuned using online experiences sampled from 40 parallelized MuJoCo environments [92], while the baselines use only one environment (matching their original implementations). Episodes terminates when they reach maximum episode lengths (shown in Table A6) or all four subtasks are completed. Detailed hyperparameters are listed in Table A8 — we follow the hyperparameter choices from the original implementations of the baselines.
Larger variance with DPPO in Fig. 6.
In Fig. 6, it is shown that DPPO exhibits a larger variance in normalized score with Kitchen-Partial-v0 than Cal-QL. This is due to DPPO solving either 3/4 or 2/4 subtasks in one seed (low variance within the evaluation episodes in one seed) but high variance over seeds, whereas Cal-QL has higher variance among evaluation episodes in one seed but on average over seeds it shows lower variance. This also highlights a notable property of DPPO: Kitchen-Partial-v0 and Kitchen-Mixed-v0 have trajectories only completing subtasks, thus being highly multi-modal. Diffusion policy can sometimes struggle to learn all the modes from pre-training, and since DPPO directly fine-tunes the pre-trained policy, it can fail to converge to 100% success rate at fine-tuning. Cal-QL instead learns from all offline data during fine-tuning in an off-policy manner, thus less sensitive to pre-training performance. Nonetheless, with offline data completing tasks consistently despite varying quality (Robomimic and Kitchen-Complete-v0, which, we believe, are more realistic in the current paradigm of robot manipulation), DPPO demonstrates much better final performance than Cal-QL and other baselines in Fig. 6.
E.6 Additional details of Robomimic tasks and training in Section 5.3
@@
E.7 Descriptions of policy parameterization baselines in Section 5.3
@@
We consider unimodal Gaussian with diagonal covariance, the most commonly used policy parameterization in RL. The standard deviation for each action dimension, $\sigma_{\text{Gau}}$, is fixed during pre-training; we also tried to learn $\sigma_{\text{Gau}}$ from the dataset but we find the training very unstable. During fine-tuning $\sigma_{\text{Gau}}$ is learned starting from the same fixed value and also clipped between $0.01$ and $0.2$. Additionally we clip the sampled action to be within $3$ standard deviation from the mean. As discusses in Appendix B, we choose the PPO clipping ratio $\varepsilon$ based on the empirical clipping fraction in each task. This setup is also used in the Furniture-Bench experiments. We note that we spend significant amount of efforts tuning the Gaussian baseline, and our results with it are some of the best known ones in RL training for long-horizon manipulation tasks (exceeding our initial expectations), e.g., reaching $\sim$100success rate in Lamp with Low randomness.
@@
E.8 Additional details of Furniture-Bench tasks and training in Section 5.4
@@
To deploy state-based policies on real hardware, we utilize AprilTags [98] for part pose estimation. The Furniture-Bench [36] task suite provides AprilTags for each part and code for estimating part poses from tag detections. The process involves several steps: (1) detecting tags in the camera frame, (2) mapping tag detections to the robot frame for policy compatibility, (3) utilizing known offsets between tags and object centers in the simulator, and (4) calibrating the camera pose using an AprilTag at a known position relative to the robot base. Despite general accuracy, detections can be noisy, especially during movement or partial occlusion, which the One-leg task features. Since the task requires high precision, we find the following to help make the estimation reliable enough:
  • Camera coverage: We find detection quality sensitive to distance and angle between the camera and tag. This issue is likely due to the RealSense D435 camera having mediocre image quality and clarity and the relatively small tags. To remedy this, we opt to use 4 cameras roughly evenly spread out around the scene to ensure that at least one camera has a solid view of a tag on all the parts (i.e., as close as possible with a straight-on view). To find the best camera positions, we start with having a camera in each of the cardinal directions around the scene. Then, we adjust the pose of each to get it as close as possible to the objects while still covering the necessary workspace and capturing the base tag for calibration. Moving the robot arm around the scene to avoid the worst occlusion is also helpful.
  • Lighting: Even with better camera coverage and placement, detection quality depends on having crisp images. We find proper lighting helpful to improve image quality. In particular, the scene should be well and evenly lit around the scene without causing reflections in either the tag or table.
  • Filtering: Bad detections can sometimes cause the resulting pose estimate to deviate significantly from the true pose, i.e., jumping several centimeters from one frame to the next. This usually only happens on isolated frames, and thus before “accepting” a given detection, we check if the new position and orientation are within 5 cm and 20 degrees of the previously accepted pose. In addition, we apply low-pass filtering on the detection using a simple exponential average (with $\alpha=0.25$) to smooth out the high-frequency noise.
  • Averaging: The objects have multiple tags that can be detected from multiple cameras. After performing the filtering step, we average all pose estimates for the same object across different tags and cameras, which also helps smooth out noise. This alone, however, does not fully cancel the case when a single detection has a large jump, as this can severely skew the average, still necessitating a filtering step. Having multiple cameras benefits this step, too, as it provides more detections to average over.
  • Caching part pose in hand: A particularly difficult phase of the task to achieve good detections is when the robot transports the table leg from the initial position to the tabletop for insertion. The main problems are that the movement can blur the images, and the grasping can cause occlusions. Therefore, we found it helpful to assume that once the part was grasped by the robot, it would not move in the grasp until the gripper opened. With this, we can “cache” the pose of the part relative to the end-effector once the object is fully grasped and use this instead of relying on detections during the movement.
  • Normalization pitfalls and clipping: We generally use min-max normalization of the state observations to ensure observations are in $[-1,1]$. The tabletop part moves very little in the $z$-direction demonstration data, meaning the resulting normalization limits (the minimum and maximum value of the data) can be very close, $x_{\text{max}}-x_{\text{min}}\approx 0$. With these tight limits, the noise in the real-world detection can be amplified greatly as $x_{\text{norm}}=\frac{x-x_{\text{min}}}{x_{\text{max}}-x_{\text{min}}}$. Therefore, ensure that normalization ranges are reasonable. As an extra safeguard, clipping the data to $[-1,1]$ can also help.
  • Only estimate necessary states: Despite the One-leg task having 5 parts, only 2 are manipulated. Only estimating the pose of those parts can eliminate a lot of noise. In particular, the pose of the 3 legs that are not used and the obstacle (the U-shaped fixture) can be set to an arbitrary value from the dataset.
  • Visualization for debugging: We use the visualization tool MeshCat [^11] https://github.com/meshcat-dev/meshcat extensively for debugging of state estimation. The tool allows for easy visualizations of poses of all relevant objects in the scene, like the robot end-effector and parts, which makes sanity-checking the implementation far easier than looking at raw numbers.
@@
E.9 Additional details of Avoid task from D3IL and training in Section 6
@@
E.10 Listed training hyperparameters
@@
Actor learning rate 1e-4 for DPPO and 1e-5 for others (tuned from 1e-4 to 1e-5)
@@
$N_{\phi}$ 16
Buffer size 200000 120000 120000 120000
Batch size 1000
@@
Buffer size 1000000
Batch size 1000
@@
$N_{\theta}$ 128
$N_{\phi}$ 128
Buffer size 1000000 250000 250000 250000
Batch size 1000
DQL $\alpha_{\text{DQL}}$ 1
@@
Buffer eize 1000000
Batch size 1000
QSM $\alpha_{\text{QSM}}$ 10
$N_{\theta}$ 16
$N_{\phi}$ 16
Buffer size 1000000 250000 250000 250000
Batch size 1000
@@
Batch size 50000 7500 10000 10000
$K^{\prime}$ 10
Table A7: Fine-tuning hyperparameters for OpenAI Gym and Robomimic tasks when comparing diffusion-based RL methods. We list shared hyperparameters and then method-specific ones.
Task(s)
Method Parameter HalfCheetah-v2 Kitchen-Complete-v0 Kitchen-Partial-v0 Kitchen-Mixed-v0
Common $\gamma_{\textsc{env}}$ 0.99
RLPD $T_{p}$ 1
$T_{a}$ 1
$N_{\phi}$ 2 3 10310 10
$N_{\text{critic}}$ 10 5 5 5
Batch size 256
Cal-QL $T_{p}$ 1
$T_{a}$ 1
$\beta_{\text{cql}}$ 5
Batch size 256
IBRL $T_{p}$ 1
$T_{a}$ 1
$N_{\phi}$ 5
$N_{\text{critic}}$ 5
Batch size 256
DPPO $T_{p}$ 1 4 4 4
$T_{a}$ 1 4 4 4
$\sigma^{\text{exp}}_{\text{min}}$ 0.1
$\sigma^{\text{prob}}_{\text{min}}$ 0.1
$\gamma_{\textsc{denoise}}$ 0.99
GAE $\lambda$ 0.95
$N_{\theta}$ 5 10 10 10
$N_{\phi}$ 5 10 10 10
$\varepsilon$ 0.01
Batch size 10000 5600 5600 5600
$K$ 20
$K^{\prime}$ 10
Method Parameter Can, PH Square, PH Can, MH Square, MH
Common $\gamma_{\textsc{env}}$ 0.999
$T_{a}$ 1
$T_{a}$ 1
RLPD $N_{\phi}$ 3
$N_{\text{critic}}$ 5
Batch size 256
Cal-QL $\beta_{\text{cql}}$ 5
Batch size 256
IBRL $N_{\phi}$ 3
$N_{\text{critic}}$ 5
Batch size 256
DPPO $\sigma^{\text{exp}}_{\text{min}}$ 0.1
$\sigma^{\text{prob}}_{\text{min}}$ 0.1
$\gamma_{\textsc{denoise}}$ 0.9 0.9 0.99 0.99
GAE $\lambda$ 0.95
$N_{\theta}$ 10
$N_{\phi}$ 10
$\varepsilon$ 0.01
Batch size 6000 15000 8000 20000
$K$ 20
$K^{\prime}$ 10
Table A8: Fine-tuning hyperparameters for HalfCheetah-v2, Franka-Kitchen, Can, and Square (PH or MH datasets) when comparing demo-augmented RL methods. We list hyperparameters shared by all methods first, and then method-specific ones.